A series R-L-C circuit is connected to a 110 Hz ac source that has Vrms = 80.0 V . The circuit has a resistance of 65.0 Ω and an impedance of 104 Ω at this frequency.

What average power is delivered to the circuit by the source?
Answer #1

Using V = IZ

80 = I (104)

I = 0.7692 amp

Average power delivered = V×I× cos\phi

= 80(.7692)R/Z

= 61.53× 65/104

= 38.46 Watt